6'4" and 8'1"
how do i express the length in feet using mixed fractions AND improper fractions??

Answer by sarah_adam(201)

lets see the conversion of inch to ft
we know 1 ft = 12 inches
therfore 1 inch = 1/12 ft
6'4" implies 6 ft 4 inches
as 1 inch = 1/12 ft therfore 4 inch = 4/12 ft that is 1/3 ft
therfore 6'4" as a mixed fraction in terms of ft would be 6 1/3 ft
the same which in improper fraction would be 19/2 ft
for 6'4"
mixed fraction : 6 1/3 ft
improper fraction : 19/3 ft (6*3 + 1)/3
8'1" implies 8 ft 1inches
as 1 inch = 1/12 ft
therfore 8'1" as a mixed fraction in terms of ft would be 8 1/12 ft
the same which in improper fraction would be 97/12 ft
for 8'1"
mixed fraction : 8 1/12 ft
improper fraction : 97/12 ft (6*3 + 1)/3